OpenStudy (anonymous):

amiright? integrate: ((e^(2x)) / e^x cant i treat that at ((e^(2x)) * e^(-x) which would just simplify to e^x???

OpenStudy (anonymous):

and even I can integrate e^x

OpenStudy (anonymous):

hehe o please help...i'm dying here

OpenStudy (anonymous):

yes , you are completely right because we subtract powers in division , and don't forget to write C (constant of integration) I often forget it :(

OpenStudy (turingtest):

yes and yes

OpenStudy (anonymous):

yay me...that question has stumped me for a week

OpenStudy (turingtest):

\[\int e^xdx=e^x+C\]\[\frac d{dx}e^x=e^x\]so those are pretty easy to remember ;)

OpenStudy (anonymous):

luckily no need for c the real question is does that converge or diverge and since e^x heads for infinity I can say diverge with more than my usual 50/50 chance of getting it right thanks all

OpenStudy (turingtest):

if it's\[\int_{a}^{\infty}e^xdx\]or any integral over +/-infinity of e^x like that, yeah, it diverges

OpenStudy (turingtest):

sorry, NOT if it's negative infinity

OpenStudy (turingtest):

\[\int_{-\infty}^{a}e^xdx\]converges

OpenStudy (anonymous):

ah good point - i will keep an eye out for that on the test tomorrow...all of our practice questions have been positive though so that would be cruel

OpenStudy (turingtest):

good luck!

OpenStudy (anonymous):

good luck with your Exam :)